# Approaching ballistic motion in 3D simulations of gamma-ray burst jets in realistic binary neutron star merger environments

### v1

@@ 3.2 Energy budget and ballistic regime
In this Section, we discuss the evolution of the jet energetics with time, including also the data for the first seconds of evolution from the reference simulation.
The upper panel of Fig. 6 shows the evolution of the kinetic ($E_{\mathrm{kin}}$), internal ($E_{\mathrm{int}}$), and magnetic ($E_{\mathrm{mag}}$) energies over the whole domain. In the very first milliseconds of evolution all the energetic components decrease, due to the effects of the gravitational attraction from the inner radial boundary (see P23). Subsequently, internal and magnetic components are continuously converted into kinetic form.
Focusing the attention on the extended evolution ($t\!>\!t_{\mathrm{jet}}+3$ s), we observe that the energy $E_{\mathrm{sum}}\equiv E_{\mathrm{kin}}+E_{\mathrm{int}}+E_{\mathrm{mag}}$ in the new Cartesian domain is approximately $4.1\times 10^{50}$ erg, very well conserved throughout the extended simulation. At the beginning, energy conservation is slightly altered by the spurious injection of material at the lower $y$ boundary. This effect is however removed after the first second, due to the exponential decay imposed on $v_{y}$ at that boundary (see above). Quantitatively, $E_{\mathrm{sum}}$ increases by about 0.7% in the first second (from 3 to 4 s after jet launching) and by only 0.08% from 4 s to the end. Such a good level of conservation shows that the increase of kinetic energy corresponds precisely to the decrease in magnetic and internal energies and therefore no significant errors are present in the numerical simulation nor in the way the energies are computed over the computational domain.
@@
Figure 6: Upper panel: Time evolution of the different energy components in the whole computational domain: kinetic ($E_{\mathrm{kin}}$), internal ($E_{\mathrm{int}}$), and magnetic ($E_{\mathrm{mag}}$) energies are represented, with dashed (continuous) lines referring to the evolution employing the spherical (Cartesian) grid. The black vertical line represents the breakout time (see text). Lower panel: Evolution of the fraction ${E}_{\mathrm{kin}}/{E}_{\mathrm{sum}}$ (see text) within the simulation time in the original spherical grid (dashed blue line) and after remapping into the new Cartesian grid (full blue line), further extrapolated to later times via a polynomial fit (orange line). Time is represented in logarithmic scale. In the inset, we show with a linear scale in time the behaviour of the same energy ratio referred only to the extended Cartesian simulation.
@@ 4 Jet structure
Here, we analyse the main properties of the most energetic part of the jet, quantitatively defined as ‘head’ (see 2.3). Specifically, we focus on two of the relevant quantities for afterglow emission, namely, the radial and angular distributions of energy and Lorentz factor.
Hereafter, the relevant data are remapped in spherical coordinates ($r,\theta,\phi$), as described in 2.3.
@@ 4.1 Energy distribution
Figure 9: 2D angular distribution of the isotropic equivalent energy (top) and the radially averaged Lorentz factor (bottom) of the jet head at times $t_{BO}$, $t_{1}$, $t_{2}$ and $t_{3}$. The jet injection axis corresponds to the central point in each panel.
@@ 4.2 Front-view angular distributions
In this Section, we discuss the angular distribution of Lorentz factor and energy per unit solid angle characterizing the jet head.
In order to have a representative value of the Lorentz factor for each angular position with respect to the jet injection axis, we define an energy-weighted radial average $\bar{\Gamma}$ as follows:
@@
where $(\Theta,\Phi)$ indicate the polar and azimuthal angles with respect to the jet injection axis, (not to be confused with the actual polar coordinates $(\theta,\phi)$, defined with respect to a polar axis orthogonal to the injection axis, see Sect. 2). The choice of weighting over $e_{\mathrm{sum}}$ is motivated by the fact that this represents the available energy which could be converted into radiation.
Next, we consider the distribution of the isotropic equivalent energy. In our case it is defined as:
@@
The overall shape of the jet remains mostly unchanged throughout the Cartesian simulation, consistently with the fact that the jet is not anymore interacting with the post-merger environment. However, as shown by the quantitative analysis which will be presented in the next Section and further discussed in Sect. 5.1, there are still differences between $t_{1}$ and $t_{3}$ that are significant for predicting the afterglow emission.
@@ 4.3 Angular profiles
In the following analysis, the angular profiles of $\bar{\Gamma}(\Theta,\Phi)$ and $E_{\mathrm{iso}}(\Theta,\Phi)$ (see Sect. 4.2) are examined as functions of $\Theta$ alone for each selected value of the jet azimuthal angle $\Phi$ [^4] The impact of the spherical-to-Cartesian and Cartesian-to-spherical interpolations (Sect. 2.2 and 2.3) on the accuracy of such angular profiles is addressed in Appendix A.1.
The resulting profiles are presented in Fig. 10, where in each panel a number of lines referring to different values of $\Phi$ are shown together with the profile obtained by averaging over $\Phi$ (thick line), namely:
@@
Figure 11: Analogously to Fig. 10, we show here the $\Phi$-averaged profiles of the radially averaged Lorentz factor (left) and isotropic equivalent energy (right) of the jet head as functions of $\Theta$. In order to highlight how the angular distributions evolve with time, in each panel we plot together the result for four different times.
@@ 5.1 Saturation limit
For an outflow with uniform Lorentz factor $\Gamma$ and kinetic energy
$E_{\mathrm{kin}}=\Gamma(\Gamma-1)m_{0}c^{2}\,,$
where $m_{0}$ is a representative mass, we can compute a saturation limit by substituting $E_{\mathrm{kin}}$ with the total available energy $E_{\mathrm{sum}}$, thus assuming that all the energy has been converted into kinetic:
@@
Based on the above, multiplying by factor $\sqrt{E_{\mathrm{sum}}/E_{\mathrm{kin}}}$ our $\bar{\Gamma}(\Theta)$ at a given time would give a general estimate of the corresponding saturation profile $\bar{\Gamma}_{s}(\Theta)$. A more precise estimate can be obtained by considering the square root of the local ratio of total and kinetic energy densities $\sqrt{e_{\mathrm{sum}}/e_{\mathrm{kin}}}$, to be applied for individual fluid elements. We thus compute the saturation radially averaged Lorentz factor of the jet head $\bar{\Gamma}_{s}$ as (see Eq. (2))
@@
and then obtain the $\Phi$-averaged profile $\bar{\Gamma}_{s}(\Theta)$.
@@
Then, comparing the two saturation profiles at 3 s and 9 s, the discrepancy is only about 2% at the peak, but it increases to 8% (35%) at smaller (larger) $\Theta$. Since afterglow emission is rather sensitive to the input Lorentz factor profile, the overall light curve may be significantly different if predicted using the information at 3 s or at 9 s. We will quantify such differences in a future work, where the afterglow emission will be investigated for a set of jet simulations including the one considered here.
As shown by the above analysis, extending a 3D sGRB jet simulation up to several seconds or more can be useful to assess the inevitable uncertainties on the afterglow emission related to the residual jet evolution.
@@ 5.2 Deceleration radius
Considering a relativistic fireball travelling in a uniform ISM with baryon number density $n$, the deceleration radius is defined by the condition that the isotropic-equivalent mass of the ISM swept-up by the jet equals its initial energy $E_{\mathrm{iso}}$ (Rees & Meszaros, 1992; Salafia & Ghirlanda, 2022), namely:
@@
In Fig. 13 we show the results for $R_{\mathrm{dec}}$ obtained from the $E_{\mathrm{iso}}$ and $\bar{\Gamma}$ distributions at the end of the simulation, assuming $n=10^{-3}$ cm-3. The upper plot represents a front-view angular distribution, while the lower panel is obtained by using the azimuthally averaged profiles, showing the deceleration radius as a function of $\Theta$. Since the azimuthal symmetry assumption is most used in literature, we focus the following analysis on the lower panel. We find $R_{\mathrm{dec}}(\Theta)$ $\sim[2.7-4.3]\times 10^{18}$ cm, barely consistent with the typical range $10^{16}-10^{18}$ cm for (long) GRBs (Mészáros, 2006; Meszaros & Rees, 1993a). Typical values of $n$ for short GRBs have been constrained in the ample range $[10^{-5}-1]$ cm-3 (e.g., Salafia & Ghirlanda 2022), thus allowing for values of $R_{\mathrm{dec}}$ from 10 times smaller to a few times larger.
@@ 6 Summary and discussion
We remark that, for the injection parameters employed here, the jet has to spend a large fraction of its energy to drill through the post-merger environment, with a final Lorentz factor limited to slightly more than 10. While such a jet is unlikely powerful enough to produce a canonical sGRB, it still represents a possible physical case. The systematic application to a number of different jet injection parameters (e.g. magnetization, initial luminosity, among others), including cases corresponding to typical sGRB jets, will be the subject of future work. We further plan to compute afterglow emission light curves for each case and compare them with the observations of GRB 170817A to plausibly obtain constraints on the physical conditions at the jet injection.
@@ A.1 Interpolation accuracy
Here we quantify the impact of interpolating the simulation data from a spherical to a Cartesian grid (Sect. 2.2), and back to a new spherical grid (Sect. 2.3), which is necessary to analyse the outcome of the evolution in terms of angular distributions.
In particular, we compare $\bar{\Gamma}(\Theta)$ and $E_{\mathrm{iso}}(\Theta)$ at $t-t_{\mathrm{jet}}=3$ s obtained (i) from the original evolution in spherical coordinates and (ii) after applying a double interpolation spherical-to-Cartesian and Cartesian-to-spherical as employed in our work. We note that the resolution of the Cartesian grid is slightly lower than the spherical one closer to the origin, and higher elsewhere (including within the head of the jet).
The comparisons are shown in Fig. LABEL:fig:domcfr, for both $\bar{\Gamma}$ (top panel) and $E_{\mathrm{iso}}$ (bottom panel). Four different profiles are considered (in different colours), referring to specific values of the azimuthal angle $\Phi$ and the $\Phi$-averaged result. The resulting profiles are slightly different, as expected after a double interpolation changing coordinates and resolution, but still in rather good accordance. We find a maximum relative difference $\lesssim 10$ for both quantities.
We remark that in the passage from the original spherical grid to the Cartesian one, which is used to continue the evolution after 3 s from the jet launching time, there is no loss of resolution affecting the results of the simulation.
@@ A.2 Jet head radial extension
We investigate the solidity of our definition for the jet head. Figure LABEL:fig:treshcfr shows the $\Phi$-averaged profiles of $\bar{\Gamma}$ and $E_{\mathrm{iso}}$ at 3.5 s after jet launching time, varying the threshold fraction of the maximum $dE/dt$ adopted to define the radial extension of the jet head (see Sect. 2.3). For both quantities, the profiles show that using a threshold of 10, 15, or 20% gives very similar quantitative results, thus justifying our choice of 15% as fiducial fraction.

### v2

@@ 3.2 Energy budget and ballistic regime
In this Section, we discuss the evolution of the jet energetics with time, including also the data for the first seconds of evolution from the reference simulation. The upper panel of Fig. 6 shows the evolution of the kinetic ($E_{\mathrm{kin}}$), internal ($E_{\mathrm{int}}$), and magnetic ($E_{\mathrm{mag}}$) energies over the whole domain. In the very first milliseconds of evolution all the energetic components decrease, due to the effects of the gravitational attraction from the inner radial boundary (see P23). Subsequently, internal and magnetic components are continuously converted into kinetic form.
Focusing the attention on the extended evolution ($t\!>\!t_{\mathrm{jet}}+3$ s), we observe that the energy $E_{\mathrm{sum}}\equiv E_{\mathrm{kin}}+E_{\mathrm{int}}+E_{\mathrm{mag}}$ in the new Cartesian domain is approximately $4.1\times 10^{50}$ erg, very well conserved throughout the extended simulation. At the beginning, energy conservation is slightly altered by the spurious injection of material at the lower $y$ boundary. However, this effect is removed after the first second due to the exponential decay imposed on $v_{y}$ at that boundary (see above). Quantitatively, $E_{\mathrm{sum}}$ increases by about 0.7% in the first second (from 3 to 4 s after jet launching) and by only 0.08% from 4 s to the end. Such a good level of conservation shows that the increase of kinetic energy corresponds precisely to the decrease in magnetic and internal energies, and therefore no significant errors are present in the numerical simulation nor in the way the energies were computed over the computational domain.
@@
Figure 6: Time evolution of the different energy components in the whole computational domain. Upper panel: kinetic ($E_{\mathrm{kin}}$), internal ($E_{\mathrm{int}}$), and magnetic ($E_{\mathrm{mag}}$) energies are represented, with dashed (continuous) lines referring to the evolution employing the spherical (Cartesian) grid. The black vertical line represents the breakout time (see text). Lower panel: Evolution of the fraction ${E}_{\mathrm{kin}}/{E}_{\mathrm{sum}}$ (see text) within the simulation time in the original spherical grid (dashed blue line) and after remapping into the new Cartesian grid (full blue line), further extrapolated to later times via a polynomial fit (orange line). Time is represented in logarithmic scale. In the inset, we show with a linear scale in time the behaviour of the same energy ratio referred only to the extended Cartesian simulation.
@@ 4 Jet structure
Here, we analyse the main properties of the most energetic part of the jet, quantitatively defined as ‘head’ (see 2.3). Specifically, we focus on two of the relevant quantities for afterglow emission, namely, the radial and angular distributions of energy and Lorentz factor. Hereafter, the relevant data are remapped in spherical coordinates ($r,\theta,\phi$), as described in 2.3.
@@ 4.1 Energy distribution
Figure 9: Two-dimensional angular distribution of the isotropic equivalent energy (top) and the radially averaged Lorentz factor (bottom) of the jet head at times $t_{BO}$, $t_{1}$, $t_{2}$ and $t_{3}$. The jet injection axis corresponds to the central point in each panel.
@@ 4.2 Front-view angular distributions
In this Section, we discuss the angular distribution of Lorentz factor and energy per unit solid angle characterising the jet head. In order to have a representative value of the Lorentz factor for each angular position with respect to the jet injection axis, we define an energy-weighted radial average $\bar{\Gamma}$ as follows:
@@
where $(\Theta,\Phi)$ indicate the polar and azimuthal angles with respect to the jet injection axis, (not to be confused with the actual polar coordinates $(\theta,\phi)$, defined with respect to a polar axis orthogonal to the injection axis, see Sect. 2). The choice of weighting over $e_{\mathrm{sum}}$ is motivated by the fact that this represents the available energy that could be converted into radiation.
Next, we consider the distribution of the isotropic equivalent energy. In our case it is defined as
@@
The overall shape of the jet remains mostly unchanged throughout the Cartesian simulation, consistently with the fact that the jet is not anymore interacting with the post-merger environment. However, as shown by the quantitative analysis presented in the next Section and further discussed in Sect. 5.1, there are still differences between $t_{1}$ and $t_{3}$ that are significant for predicting the afterglow emission.
@@ 4.3 Angular profiles
In the following analysis, the angular profiles of $\bar{\Gamma}(\Theta,\Phi)$ and $E_{\mathrm{iso}}(\Theta,\Phi)$ (see Sect. 4.2) are examined as functions of $\Theta$ alone for each selected value of the jet azimuthal angle $\Phi$ [^4] The impact of the spherical-to-Cartesian and Cartesian-to-spherical interpolations (Sect. 2.2 and 2.3) on the accuracy of such angular profiles is addressed in Appendix A.1. The resulting profiles are presented in Fig. 10, where in each panel a number of lines referring to different values of $\Phi$ are shown together with the profile obtained by averaging over $\Phi$ (thick line), namely:
@@
Figure 11: $\Phi$-averaged profiles (see Fig. 10) of the radially averaged Lorentz factor (left) and isotropic equivalent energy (right) of the jet head as functions of $\Theta$. In order to highlight how the angular distributions evolve with time, in each panel we plot together the result for four different times.
@@ 5.1 Saturation limit
The lab-frame kinetic energy density of a representative element of the outflow is given by
$e_{\mathrm{kin}}=\rho\Gamma(\Gamma-1)c^{2}\,,$
where $\Gamma$ is its Lorentz factor and $\rho$ the rest-mass density of the element in the comoving frame. We can then estimate a saturation limit by substituting $e_{\mathrm{kin}}$ with $e_{\mathrm{sum}}$, thus assuming that all the energy has been converted into kinetic, namely:
@@
We thus computed the saturation radially averaged Lorentz factor of the jet head $\bar{\Gamma}_{s}$ as (see Eq. (2))
@@
and then obtained the $\Phi$-averaged profile $\bar{\Gamma}_{s}(\Theta)$.
@@
Then, comparing the two saturation profiles at 3 s and 9 s, the discrepancy is only about 2% at the peak, but it increases to 8% (35%) at smaller (larger) $\Theta$. Since afterglow emission is rather sensitive to the input Lorentz factor profile, the overall light curve may be significantly different if predicted using the information at 3 s or at 9 s. We will quantify such differences in a future work, where the afterglow emission will be investigated for a set of jet simulations, including the one considered here. As shown by the above analysis, extending a 3D sGRB jet simulation up to several seconds or more can be useful to assess the inevitable uncertainties on the afterglow emission related to the residual jet evolution.
@@ 5.2 Deceleration radius
Considering a relativistic fireball travelling in a uniform ISM with baryon number density $n$, the deceleration radius is defined by the condition that the isotropic-equivalent mass of the ISM swept-up by the jet equals its initial isotropic-equivalent energy $E_{\mathrm{iso}}$ (Rees & Meszaros, 1992; Salafia & Ghirlanda, 2022), namely:
@@
In Fig. 13 we show the results for $R_{\mathrm{dec}}$ obtained from the $E_{\mathrm{iso}}$ and $\bar{\Gamma}$ distributions at the end of the simulation, assuming $n=10^{-3}$ cm-3. The upper plot represents a front-view angular distribution, while the lower panel is obtained by using the azimuthally averaged profiles, showing the deceleration radius as a function of $\Theta$. Since the azimuthal symmetry assumption is most used in literature, we focus the following analysis on the lower panel. We find $R_{\mathrm{dec}}(\Theta)$ $\sim[2.7-4.3]\times 10^{18}$ cm, barely consistent with the typical range $10^{16}-10^{18}$ cm for (long) GRBs (Mészáros, 2006; Meszaros & Rees, 1993a). Typical values of $n$ for short GRBs have been constrained in the ample range $[10^{-5}-1]$ cm-3 (e.g. Salafia & Ghirlanda 2022), thus allowing for values of $R_{\mathrm{dec}}$ from ten times smaller to a few times larger.
@@ 6 Summary and discussion
We remark that, for the injection parameters employed here, the jet has to spend a large fraction of its energy to drill through the post-merger environment, with a final Lorentz factor limited to slightly more than 10. While such a jet is unlikely powerful enough to produce a canonical sGRB, it still represents a possible physical case. The systematic application to a number of different jet injection parameters (e.g. magnetisation, initial luminosity, among others), including cases corresponding to typical sGRB jets, will be the subject of future work. We further plan to compute afterglow emission light curves for each case and compare them with the observations of GRB 170817A to plausibly obtain constraints on the physical conditions at the jet injection.
@@ A.1 Interpolation accuracy
Here we quantify the impact of interpolating the simulation data from a spherical to a Cartesian grid (Sect. 2.2), and back to a new spherical grid (Sect. 2.3), which is necessary to analyse the outcome of the evolution in terms of angular distributions. In particular, we compare $\bar{\Gamma}(\Theta)$ and $E_{\mathrm{iso}}(\Theta)$ at $t-t_{\mathrm{jet}}=3$ s obtained (i) from the original evolution in spherical coordinates and (ii) after applying a double interpolation spherical-to-Cartesian and Cartesian-to-spherical as employed in our work. We note that the resolution of the Cartesian grid is slightly lower than the spherical one closer to the origin, and higher elsewhere (including within the head of the jet).
The comparisons are shown in Fig. LABEL:fig:domcfr, for both $\bar{\Gamma}$ (top panel) and $E_{\mathrm{iso}}$ (bottom panel). Four different profiles are considered (in different colours), referring to specific values of the azimuthal angle $\Phi$ and the $\Phi$-averaged result. The resulting profiles are slightly different, as expected after a double interpolation changing coordinates and resolution, but still in rather good accordance. We find a maximum relative difference $\lesssim 10$ for both quantities. We remark that in the passage from the original spherical grid to the Cartesian one, which is used to continue the evolution after 3 s from the jet launching time, there is no loss of resolution affecting the results of the simulation.
@@ A.2 Jet head radial extension
We investigate the solidity of our definition for the jet head. Figure LABEL:fig:treshcfr shows the $\Phi$-averaged profiles of $\bar{\Gamma}$ and $E_{\mathrm{iso}}$ at 3.5 s after jet launching time, varying the threshold fraction of the maximum $dE/dt$ adopted to define the radial extension of the jet head (see Sect. 2.3). For both quantities, the profiles show that using a threshold of 10%, 15%, or 20% gives very similar quantitative results, thus justifying our choice of 15% as fiducial fraction.